Does the function f(x,y)= (x^2-y^2)/(x^2+y^2), x
1
Expert's answer
2022-01-11T11:58:42-0500

Given function,


"f(x)=\dfrac{x^2-y^2}{x^2+y^2}"


let "y=vx~~~~-(1)"


"\dfrac{dy}{dx}=\dfrac{x^2-v^2x^2}{x^2+v^2x^2}"


"=\dfrac{1-v^2}{1+v^2}"


Hence The degree of numerator and denominator is same, So Given function is Homogeneous.